OpenStudy (anonymous):

what is perimeter of triangle ABC if AB=12ft and BC=16ft

OpenStudy (ikram002p):



OpenStudy (anonymous):

Use the Pythagorean theorem: a^2 + b^2 = c^2 You know that the side AB is 12 ft long, so let's replace a with 12. You know that side BC is 16 ft long, so let's replace b with 16. Now the equation becomes: 12^2 + 16^2 = c^2 once you solve for c^2, you can find the square root, which will give you the length of side CA. Okay, so now that you have all of the sides determined, add them together. 12^2 + 16^2 = c^2 144 + 256 = 400 Remember to find the square root of 400 now, because it is equivalent to c to the second power. 20*20=400, so 20 is the square root of 400 and also the length of CA. Finally, you can add the sides together, 12 + 16 + 20 = 48, and ta dum, you have your perimeter :)

OpenStudy (ikram002p):

mmmm its brety uncomplete qs since u can form many triangle with 16 & 12 but u can know the maximum perimeter that it could be . which is of the furmola max < 2 (Ab+bc) max<56 @Shirobara u cant use the Pythagorean theorem since its not mintion that its right triangle .

Directrix (directrix):

@dememinskoff Please post whether or not the given triangle is supposed to be a right triangle. Thanks.